OpenStudy (anonymous):

Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

for f(x). you may use below identity \(a^2-b^2 = (a+b)(a-b)\)

ganeshie8 (ganeshie8):

you can define f(x) like below :- \(f(x) = x^2-2^2\)

ganeshie8 (ganeshie8):

and factor it using above identity

OpenStudy (anonymous):

so we take f(x) = x^2 - 2^2 and factor it?

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

Can you walk me through factoring it?

ganeshie8 (ganeshie8):

did u look at the identity ?

ganeshie8 (ganeshie8):

my first reply...

OpenStudy (anonymous):

alright so we factor it using a^2 - b^2 = (a + b)(a - b) f(x) = x^2 - 2^2 so... x^2 - 2^2 = (x + (-2)(x - (-2)?

ganeshie8 (ganeshie8):

x^2 - 2^2 = (x + 2)(x - 2) thats it

OpenStudy (anonymous):

So thats the answer for part 1?

OpenStudy (anonymous):

you there @ganeshie8

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

for g(x), just create a quadratic wid SUM of two squares :- \(g(x) = x^2+2^2\)

ganeshie8 (ganeshie8):

you cannot factor this, so nothing much to do. move to h(x)

ganeshie8 (ganeshie8):

i want you try h(x), can u ? :)

OpenStudy (anonymous):

one sec this takes time lol

OpenStudy (anonymous):

(x − 5) * (x − 5)?

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

:D

ganeshie8 (ganeshie8):

multiply and show it as trinomial like :- \(ax^2+bx+c\)

ganeshie8 (ganeshie8):

\(h(x) = (x-5)^2 = ?\)

OpenStudy (anonymous):

I'm not sure

ganeshie8 (ganeshie8):

\(h(x) = (x-5)^2 = (x-5)(x-5) = ?\)

ganeshie8 (ganeshie8):

you knw how to multiply right ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

-5x^2? lol

ganeshie8 (ganeshie8):

yo wat happened to u.. u use to be so good few days back on these :o

ganeshie8 (ganeshie8):

\((x-5)(x-5)\) FOIL \((x^2-5x - 5x +25\) \(x^2-10x +25\) \(h(x) = x^2-10x+25\)

ganeshie8 (ganeshie8):

Fine ?

OpenStudy (anonymous):

lol been busy

ganeshie8 (ganeshie8):

I can see, you need to revise quick and get back on track ok

ganeshie8 (ganeshie8):

you should knw how to multiply (x-5)(x-5)

OpenStudy (anonymous):

alright I think I'm good, I forgot about foil lol

ganeshie8 (ganeshie8):

ive see u doing these before.. .just now you're acting like u lost all ur memory

ganeshie8 (ganeshie8):

*seen

ganeshie8 (ganeshie8):

for j(x), try something like below :- \(j(x) = 2x^2 + 4\)

ganeshie8 (ganeshie8):

you can oly factor out the GCF there, wats the GCF above ?

OpenStudy (anonymous):

2

ganeshie8 (ganeshie8):

good :) factor it out

ganeshie8 (ganeshie8):

\(j(x) = 2x^2 + 4 = 2(x^2+2)\)

ganeshie8 (ganeshie8):

thats it ! we have 4 functions, and factored them also

OpenStudy (anonymous):

Thanks man! Sorry I'm a bit rusty, been a while since I've delt with factoring xD

ganeshie8 (ganeshie8):

The function f(x) is a difference of squares. \(f(x) = x^2-4\) factored form : \(f(x) = (x+2)(x-2)\) The function g(x) is a sum of squares. \(f(x) = x^2+4\) factored form : \(CANNOT ~ FACTOR\) The function h(x) is a perfect square trinomial. \(f(x) = x^2-10x+25\) factored form : \(f(x) = (x-5)(x-5)\) The function j(x) can only have a GCF factored out of it. \(f(x) = 2x^2+4\) factored form : \(f(x) = 2(x^2+2)\)

ganeshie8 (ganeshie8):

Enjoy...

ganeshie8 (ganeshie8):

hey np, it wont take much time for u to revise things...